AMS-TEX offers tfrac, frac, and dfrac, all useful.

fg = \frac{1}{2}[(f + g)^{2} − f^{2} − g^{2}]

\frac{x}{1 + \frac{x}{2}} versus \frac{x}{1 + \frac{x}{2}}

AMS-TEX also offers cfrac for continued fractions.

a_{0} + 1\frac{1}{a_{1} + \frac{1}{a_{2} + \frac{1}{a_{3} + \frac{1}{a_{4}}}}}